OpenStudy (anonymous):

the base of a triangle is 3cm longer than its height. If the area of the triangle is 27cm^2 what is its height?

OpenStudy (anonymous):

The area of a triangle is 1/2 *base*height. A=0.5bh b=3 A=27 Now we can solve for h.

OpenStudy (anonymous):

A = (1/2) bh h = x b = ? What is the base?

OpenStudy (saifoo.khan):

B = air base.

OpenStudy (anonymous):

air base?